OpenStudy (anonymous):

How many times more intense is the sound of a space shuttle launch (180dB) than the sound of jet engine (140 dB)?

OpenStudy (anonymous):

180/140=1.28 times

OpenStudy (anonymous):

did you use this formula: L=10log(I/I0)?

OpenStudy (anonymous):

i guess no

OpenStudy (anonymous):

I'm suppose to..

OpenStudy (anonymous):

but it's just ratio...of two quantities

OpenStudy (anonymous):

sigh. Okay.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

140 = 10log(I/I0) 14=log(I/10) \[\huge 10^{14} =10^{\log_{ } I/10}\] \[\huge 10^{14} = I/10\] \[\huge 10^{15} =I\]

OpenStudy (anonymous):

find I for 180dB divide by 10^15